OpenStudy (anonymous):

find roots: x^4+2x^3+3x^2+2x+1=0

OpenStudy (anonymous):

First of all divide the whole equation with x^2 to make it quadratic equation then we will we the roots of this equation.

OpenStudy (anonymous):

but if ewe do it then last two terms have -ve exponents.then?

OpenStudy (anonymous):

Ntice that \[ x^4 + 2 x^3 + 3 x^2 + 2 x + 1 == (1 + x + x^2)^2 \] It should be easy now to use the quadratic formula.

OpenStudy (anonymous):

To otain the above notice that \[ x^4+2 x^3+3 x^2+2 x+1=\\x^4+2 x^3+2 x^2+x^2+2 x+1=\\x^4+2 (x+1) x^2+(x+1)^2=\left(x^2+x+1\right)^2\\ \]